# Electron charge coherence on a solid neon surface

### v1

@@ I Introduction
Helium (He) and neon (Ne), noble-gas elements with extreme chemical inertia, spontaneously condense into ultrapure quantum liquids and solids at low temperature and pressure [13]. Compared with classical liquids and solids, condensed He and Ne have negligible charged excitations or spinful isotopes (3He and 21Ne). They offer a vacuum-like environment with minimal charge and spin noises for electron qubits [14, 13]. When an excess electron approaches a liquid He or solid Ne from a vacuum, it can form surface states by two effects. First, there is a repulsive barrier of approximately 1 eV resulting from the Pauli exclusion between the excess electron and the shell electrons in He and Ne atoms. Second, there is an attractive well from the polarization-induced image charge inside the liquid or solid [15, 16, 17]. A surface electron can be further trapped in a quantum-dot structure by on-chip electrodes and coupled with microwave photons in a superconducting quantum circuit [18, 19, 20].
@@
In the past few years, we successfully realized the electron-on-solid-neon (eNe) motional (charge) qubits [21] and obtained charge coherence times up to 0.1 ms [22]. Despite already exceeding the coherence times of traditional superconducting and semiconductor charge qubits by several orders [9, 23], the current eNe charge coherence times are probably limited by the imperfect surface and redundant electrons trapped on it [24, 25]. With optimized growth of solid Ne and refined electron loading procedure, the eNe charge coherence times may be significantly improved.
@@
In this paper, we present a theoretical investigation of the intrinsic decoherence mechanisms of the motional (charge) states of a single electron on a perfect solid Ne surface. We calculate the corresponding relaxation and coherence times, which provide the upper limit of the eNe charge coherence times in this system. The system is illustrated in Fig. I, which mimics the thin-film structures of typical fabricated devices. A single electron hovers above the surface of a solid Ne film with a thickness denoted by $w$. The devices are typically superconducting circuits made of aluminum (Al) or niobium (Nb) film with a thickness denoted by $t$, deposited on low-microwave-loss substrates such as intrinsic silicon (Si) or sapphire (Al2O3). The solid Ne layer can be grown on top of both the metal electrode and the bare substrate.
@@
Fig. 1. : Schematic illustration of the electron-on-solid-neon (eNe) system, featuring an electron floating in the vacuum above a solid Ne film. The solid neon film can cover both the metal electrodes and the dielectric substrate of the quantum circuit device. There are longitudinal (acoustic) phonons in the bulk and their induced displacement on the surface. The electron’s wavefunction, represented by a red cloud, resides at a mean distance of $d\approx 2$ nm from the surface and extends along the surface over a scale of 100 nm. The diagram shows a harmonic ground-state wavefunction (pink curve) with a Gaussian width $a_{x}$.
II Surface electronic states
@@
In this paper, we adopt the simplest model to describe the surface electronic states [26]. We approximate the Pauli barrier to be infinitely high, which forces the electron wavefunction to vanish inside solid Ne, $z\leq 0$. We also assume the solid Ne to be thick, $w\gtrsim 100$ nm, so the influence from the metal electrode and substrate on the wavefunction is negligible. The image potential in the vertical ($z$) direction is
@@
with $e$ being the elementary charge and $\epsilon_{\scriptscriptstyle\text{Ne}}=1.244$ being the dielectric constant of uniform solid neon.
The electron’s eigenenergies are hydrogenic,
@@
They are related by $\mathit{\Lambda}=2Rr_{\scriptscriptstyle\text{B}}$. The ground-state wavefunction is
$\psi_{1}(z)=\frac{2}{\sqrt{r_{\scriptscriptstyle\text{B}}}}\frac{z}{r_{\scriptscriptstyle\text{B}}}{\rm e}^{-z/r_{\scriptscriptstyle\text{B}}}.$ (6)
At the typical experimental temperature, $T=10$ mK, the electron’s vertical motion perpendicular to the Ne surface is effectively frozen in the ground state $\left|1_{z}\right\rangle$, whereas its lateral motion parallel to the $xy$ plane is confined by on-chip electrodes. We approximate the trapping potential in each of the lateral dimensions to be harmonic, with level spacing $\hbar\omega_{x,y}$. Thus, the Hamiltonian ^$\hat{\mathcal{H}}_{\text{el}}$ for a single electron is given by
^$\hat{\mathcal{H}}_{\text{el}}=-\frac{\hbar^{2}}{2m_{e}}\nabla^{2}+V_{x}(x)+V_{y}(y)+V_{z}(z),$ (7)
@@
The in-plane motional states of the electron are denoted as $\left|n_{x},n_{y}\right\rangle$ with $n_{x},n_{y}=0,1,2,\dots$ being the quantum numbers of quantum harmonic oscillators in $x$ and $y$, respectively. Although in this model an electron’s wavefunction entirely resides in the vacuum, phonon excitations in solid Ne can still decohere electron’s emotional states, as discussed below.
@@ III Decoherence via photons
Before talking about phonons, let us first talk about photons which set the upper bound of the relaxation and coherence times in this system.
@@ III.1 Spontaneous emission in a vacuum
Here, $d_{x}=\sqrt{2}ea_{x}$ is the electric dipole moment in $x$ with the confining length $a_{x}=\left(\hbar/m\omega_{x}\right)^{1/2}$, and $f_{x}=\omega_{x}/2\pi$ is the transition frequency. The rate of radiative decay, as given by Eq. (10), determines the relaxation time $T_{1}^{\left(\text{vac}\right)}$. For our typical choice of $f_{x}=6.4$ GHz, the corresponding $T_{1}^{\left(\text{vac}\right)}$ is 999999 s. This gives a remarkably long upper bound of the coherence time $T_{2}^{\left(\text{vac}\right)}$ , ref [28],
@@ III.2 Spontaneous emission in a superconducting cavity
In practice, the electron is often situated in an on-chip superconducting resonator. The resonator alters the spontaneous emission rate due to the modified electromagnetic environment, known as the Purcell effect[29]. For a simple two-level electric dipole coupled to a single-mode resonator with the resonator frequency $f_{\text{r}}=\omega_{\text{r}}/2\pi$ through nonresonant coupling (i.e., $g\ll\left|\omega_{x}-\omega_{\text{r}}\right|$), the modified spontaneous emission rate can be described as [22]
@@
Here, $\kappa$ is the resonator linewidth (leakage rate). For example, with $g/2\pi\sim 5$ MHz and $\kappa/2\pi\sim 0.5$ MHz [22], if $\left|f_{x}-f_{\text{r}}\right|\sim 500$ MHz, the relaxation time of the electron in the cavity is shortened to $T_{1}^{\left(\text{cav}\right)}=32$ ms. As a result, the coherence time is limited with an upper bound of $T_{2}^{\left(\text{cav}\right)}\leq 64$ ms, which is much shorter than the upper bound in vacuum but still longer than the coherence times limited by other mechanisms below. It is worthwhile to note that the Purcell limited radiative decay can be suppressed by introducing Purcell filter [30] or other readout techniques.
@@ IV Decoherence via phonons
The electron’s lateral motion outside the solid Ne can couple with the longitudinal (acoustic) phonons inside solid Ne. Consequently, the electron’s charge states can undergo relaxation and dephasing via emitting phonons.
The displacement field $\bm{u}(\bm{r},t)$ of solid Ne due to the longitudinal (acoustic) phonons, in a large space of volume $V$ and undisturbed mass density $\rho_{\scriptscriptstyle\text{Ne}}=1.444$ g cm-3, can be written as
@@
where $\bm{r}=(x,y,z)$ and $\bm{q}=(q_{x},q_{y},q_{z})$ are the 3D positions and wavevectors, respectively; $a_{\bm{q}}$ and $a^{\dagger}_{\bm{q}}$ and are the phonon annihilation and creation operators of the plane-wave mode with the wavevector $\bm{q}$; $\omega_{q}=c_{\scriptscriptstyle\text{Ne}}q$ is the dispersion relation that only depends on the magnitude (not direction) of the wavevector, $q=|\bm{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}}$, $c_{\scriptscriptstyle\text{Ne}}=1.133\times 10^{5}$ cm s-1 is the longitudinal sound velocity; $\bm{q}/q$ is the unit vector in the longitudinal wavefront direction.
@@
$\bm{v}(\bm{r},t)=\sum_{\bm{q}}\frac{\bm{q}}{{\rm i}q}\sqrt{\frac{\hbar\omega_{q}}{2V\rho_{\scriptscriptstyle\text{Ne}}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}+a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}}.$ (14)
The acoustic velocity potential $\mathit{\Phi}(\bm{r},t)$ is
@@
$\partial_{t}\varrho(\bm{r},t)+\rho_{\scriptscriptstyle\text{Ne}}\nabla\cdot\bm{v}(\bm{r},t)=0,$ (18)
[…]
@@
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}),$ (19)
the form of which ensures that the density variation vanishes at the Ne-vacuum interface, $z=0$, where pressure is zero. The summation in Eq. (19) only includes $q_{z}>0$ to ensure that the modes are counted correctly. Likewise, the acoustic velocity potential is
@@
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}+\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (20)
@@
×$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (21)
@@
The total Hamiltonian consists of the electron part ^$\hat{\mathcal{H}}_{\text{el}}$ as given above, the phonon part ^$\hat{\mathcal{H}}_{\text{ph}}$, and the electron-phonon coupling part ^$\hat{\mathcal{H}}_{\text{el-ph}}$,
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\text{el}}+\hat{\mathcal{H}}_{\text{ph}}+\hat{\mathcal{H}}_{\text{el-ph}},$ (22)
@@
^$\hat{\mathcal{H}}_{\text{ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a^{\dagger}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}},$ (23)
@@
^$\hat{\mathcal{H}}_{\text{el-ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z){\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (24)
@@
The relaxation rate corresponding to electron scattered by a single phonon can be calculated from the Fermi golden rule,
$\mathit{\Gamma}$ $=\frac{2\pi}{\hbar}\left|\langle n_{\text{ph}}^{f}|\langle 0|\hat{\mathcal{H}}_{\text{el-ph}}|1\rangle|n_{\text{ph}}^{i}\rangle\right|^{2}\times\delta(E_{n_{\text{ph}}}^{f}-E_{n_{\text{ph}}}^{i}-\hbar\omega_{x}),$ (25)
where |$\ket{0}$ and |$\ket{1}$ are the electron’s ground and excited state, |$\ket{n_{p}^{i}}$ and |$\ket{n_{p}^{f}}$ are the phonon’s initial and final states and $|E_{n_{\text{ph}}}^{f}-E_{n_{\text{ph}}}^{i}|=\hbar\omega_{q}$ is the energy of creation of one phonon with the wave vector $\bm{q}$. In addition, although not written explicitly here, this expression needs to be summed up in all possible final phonon states and averaged in all thermally excited initial states.
We write the electron-phonon coupling Hamiltonian via the modulation of dielectric constant as ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$ and the displacement of neon surface as ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{dis})}$. The relaxation rate corresponding to each mechanisms and different thermal phonons can be expressed as
@@
×$\times\delta(\sum_{q}\hbar\omega_{q}n_{q}-\sum_{q}\hbar\omega_{q}n^{\prime}_{q}-\hbar\omega_{x}),$
@@
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (26)
$\mathit{\Gamma}^{(\text{mod})}$ $=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\langle n_{q}|\langle 0_{x},0_{y},1_{z}|\hat{\mathcal{H}}^{(\text{mod})}_{\text{el-ph}}|1_{x},0_{y},1_{z}\rangle|n^{\prime}_{q}\rangle\right|^{2}$
×$\times\delta(\sum_{q}\hbar\omega_{q}n_{q}-\sum_{q}\hbar\omega_{q}n^{\prime}_{q}-\hbar\omega_{x}),$
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (27)
where $|1_{z}\rangle$ represents the electron’s Rydberg ground state in $z$, $|0_{x},0_{y}\rangle$ and $|1_{x},0_{y}\rangle$ represent the electron’s in-plane ground state and first excited state in $x$. The subscript “10” denotes the excited to ground state transition and $n_{q}=1/[\exp(\hbar\omega_{q}/(k_{\scriptscriptstyle\text{B}}T))-1]\ll 1$ denotes the average thermal occupation number of the phonon state with the wavenumber $q$ at $T=10$ mK. The function $G_{\bm{q}_{\scriptscriptstyle\parallel}}$ is
@@
The phonon-induced dephasing of electrons following the treatment used in ref [32]. The ensemble-averaged mean-square phase increment is given by
[…]
@@
is the energy difference between two electronic states. The energy correlator can be approximated by a $\delta$-function which gives the depahsing rate $\Gamma_{\phi}$. Similar to the relaxation calculation, the matrix element for Eq. (31) is given by
$\bra{n_{q}}E_{10}(t^{\prime})E_{10}(t^{\prime\prime})\ket{n_{q}}$ (32)
@@
×$\times[(n_{q}+1)e^{-i\omega_{q}(t^{\prime}-t^{\prime\prime})}-n_{q}e^{i\omega_{q}(t^{\prime}-t^{\prime\prime})}].$
The integral over time can be found as
[…]
@@
$\mathit{\Gamma}_{\phi}=$ $\frac{2\pi}{\hbar^{2}c_{\text{Ne}}}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left[(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}-(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}\right]^{2}\delta(q).$ (34)
For one-phonon processes, this summation is zero due to the $\delta$-function, whereas for two-phonon processes, the $\delta$-function becomes $\delta(|q_{1}|-|q_{2}|)$, meaning the annihilation of a phonon with $q_{1}$ and creation one with $q_{2}$. As seen below, since $\mathcal{V}^{\text{(mod)}}$ does not contain $\bm{p}_{\scriptscriptstyle\parallel},\bm{r}_{\scriptscriptstyle\parallel}$, it means $\left|\mathcal{V}^{\text{(mod)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{11}=\left|\mathcal{V}^{\text{(mod)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{00}=\text{constant}$. Thus, the sum is essentially determined by $\left[({\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}-({\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}\right]^{2}=\frac{1}{4}\left(q_{x}a_{x}\right)^{4}{\rm e}^{-(q_{x}^{2}a_{x}^{2}+q_{y}^{2}a_{y}^{2})/2}$. (See supplementary materials of ref [31].) For $\mathcal{V}^{\text{(dis)}}$, only the first term involves $\bm{p}_{\scriptscriptstyle\parallel}$ and contributes to the sum.
We now analyze in detail the two electron-phonon interaction mechanisms.

### v2

@@ I Introduction
Helium (He) and neon (Ne), noble-gas elements with extreme chemical inertia, spontaneously condense into ultrapure quantum liquids and solids at low temperature and pressure [13]. Compared with classical liquids and solids, condensed He and Ne have negligible charged excitations or spinful isotopes (3He and 21Ne). They offer a vacuum-like environment with minimal charge and spin noises for electron qubits [14, 13]. When an excess electron approaches a liquid He or solid Ne surface from vacuum, it can form surface states by two effects. First, a repulsive barrier of approximately 1 eV resulting from the Pauli exclusion between the excess electron and the shell electrons in He and Ne atoms. Second, there is an attractive well from the polarization-induced image charge within the liquid or solid [15, 16, 17]. A surface electron can be trapped in a quantum-dot structure by on-chip electrodes and coupled with microwave photons in a superconducting quantum circuit [18, 19, 20].
@@
In the past few years, we successfully realized the electron-on-solid-neon (eNe) motional (charge) qubits [21] and obtained charge coherence times up to 0.1 ms [22]. Despite already exceeding the coherence times of traditional charge qubits by several orders [9, 23], the current eNe charge coherence times are probably limited by the imperfect surface and redundant electrons trapped on it [24, 25]. With optimized growth of solid Ne and a refined electron loading procedure, the eNe charge coherence times may be significantly improved.
@@
In this paper, we present a theoretical investigation of the intrinsic decoherence mechanisms of the motional (charge) states of a single electron on a perfect solid Ne surface. We calculate the corresponding relaxation and coherence times, which provide the upper limit of the eNe charge coherence times in this system. The system is illustrated in Fig. I, which mimics the thin-film structures of typical fabricated devices. A single electron hovers above the surface of a solid Ne film with a thickness denoted by $w$. The devices are typically superconducting circuits made of aluminum (Al) or niobium (Nb) film with a thickness denoted by $t$, deposited on low-microwave-loss substrates such as intrinsic silicon (Si) or sapphire (Al2O3). The solid Ne layer can be grown on the metal electrode and the bare substrate.
@@
Fig. 1. : Schematics of the electron-on-solid-neon (eNe) system, featuring an electron floating in the vacuum above a solid Ne film and electron-phonon interaction. The solid neon film can cover both the metal electrodes and the dielectric substrate of the quantum circuit device. The longitudinal (acoustic) phonons in the bulk induces displacement on the Ne surface. The electron’s wavefunction, represented by a red cloud, resides at a mean distance of $d\approx 2$ nm from the surface and extends along the surface over a scale of 100 nm. The diagram shows a harmonic ground-state wavefunction (pink curve) with a Gaussian width $a_{x}$.
II Surface charge states
@@
In this paper, we adopt the simplest model to describe the surface charge states [26]. We approximate the Pauli barrier to be infinitely high, which forces the electron wavefunction to vanish inside solid Ne, $z\leq 0$. We also assume the solid Ne to be thick, $w\gtrsim 100$ nm, so the influence from the metal electrode and substrate on the wavefunction is negligible. We first discuss the vertical bound states. The image charge potential in the vertical ($z$) direction is
@@
with $e$ being the elementary charge and $\epsilon_{\scriptscriptstyle\text{Ne}}=1.244$ being the dielectric constant of uniform solid neon [15].
The vertical state eigenenergies are hydrogenic,
@@
They are related by $\mathit{\Lambda}=2Rr_{\scriptscriptstyle\text{B}}$. The ground-state wavefunction $\left|1_{z}\right\rangle$ is given by
$\psi_{1}(z)=\langle\bm{r}|1_{z}\rangle=\frac{2}{\sqrt{r_{\scriptscriptstyle\text{B}}}}\frac{z}{r_{\scriptscriptstyle\text{B}}}{\rm e}^{-z/r_{\scriptscriptstyle\text{B}}}.$ (6)
At the typical experimental temperature, $T=10$ mK, the electron’s vertical motion perpendicular to the Ne surface is effectively frozen in the ground state, whereas its lateral motion parallel to the $xy$ plane is confined by the on-chip electrodes. We approximate the trapping potential in each of the lateral dimensions to be harmonic, with level spacing $\hbar\omega_{x,y}$. Thus, the Hamiltonian $\hat{\mathcal{H}}_{\text{el}}$ for a single electron is given by
$\hat{\mathcal{H}}_{\text{el}}=-\frac{\hbar^{2}}{2m_{e}}\nabla^{2}+V_{x}(x)+V_{y}(y)+V_{z}(z),$ (7)
@@
The in-plane motional states of the electron are denoted as $\left|n_{x},n_{y}\right\rangle$ with $n_{x},n_{y}=0,1,2,\dots$ being the quantum numbers of quantum harmonic oscillators in $x$ and $y$, respectively. Although in this model an electron’s wavefunction entirely resides in the vacuum, phonon excitations in solid Ne can still interact with the electron’s motional states, playing the role of a decoherence source, as discussed below.
@@ III Decoherence via photons
Before discussing phonon-induced decoherece, we first discuss photon-induced spontaneous emission, which sets the upper bound of the relaxation and coherence times in this system.
@@ III.1 Spontaneous emission in a vacuum
Here, $d_{x}=\sqrt{2}ea_{x}$ is the electric dipole moment in $x$ with the confining length $a_{x}=\left(\hbar/m\omega_{x}\right)^{1/2}$, and $f_{x}=\omega_{x}/2\pi$ is the transition frequency. The rate of radiative decay, as given by Eq. (10), determines the relaxation time $T_{1}^{\left(\text{vac}\right)}$. For our typical choice of $f_{x}=6.4$ GHz, the corresponding $T_{1}^{\left(\text{vac}\right)}$ is $99$ s. This gives a remarkably long upper bound of the coherence time $T_{2}^{\left(\text{vac}\right)}$ , ref [28],
@@ III.2 Spontaneous emission in a superconducting cavity
In practice, the electron usually resides in an on-chip superconducting resonator. The resonator alters the spontaneous emission rate due to the modified electromagnetic environment, known as the Purcell effect [29]. For a simple two-level charge state coupled to a single-mode resonator with the resonance frequency $f_{\text{r}}=\omega_{\text{r}}/2\pi$, in the dispersive regime (i.e., the coupling strength $g\ll\left|\omega_{x}-\omega_{\text{r}}\right|$), the modified spontaneous emission rate can be written as
@@
Here, $\kappa$ is the resonator linewidth (leakage rate). For the typical experimental settings, with $g/2\pi\sim 5$ MHz[22], $\kappa/2\pi\sim 0.5$ MHz and $\left|f_{x}-f_{\text{r}}\right|\sim 500$ MHz, the relaxation time of the electron in the cavity is shortened to $T_{1}^{\left(\text{cav}\right)}=32$ ms. As a result, the coherence time is limited by an upper bound of $T_{2}^{\left(\text{cav}\right)}\leq 64$ ms, which is much shorter than the upper bound in vacuum but still longer than the coherence times limited by other mechanisms below. In addition, the Purcell limited radiative decay can be suppressed by introducing Purcell filter [30] or other readout techniques.
@@ IV Decoherence via phonons
The electron’s lateral motion can couple with the longitudinal (acoustic) phonons inside solid Ne. Consequently, the electron’s charge state can undergo relaxation and dephasing via emitting phonons.
The displacement field $\bm{u}(\bm{r},t)$ of solid Ne due to the longitudinal phonons, in a large space of volume $V$ and undisturbed mass density $\rho_{\scriptscriptstyle\text{Ne}}=1.444$ g cm-3 (ref. [31]), can be written as
@@
where $\bm{r}=(x,y,z)$ and $\bm{q}=(q_{x},q_{y},q_{z})$ are the 3D positions and wavevectors, respectively; $a_{\bm{q}}$ and $a^{\dagger}_{\bm{q}}$ and are the phonon annihilation and creation operators of the plane-wave mode with the wavevector $\bm{q}$; $\omega_{q}=c_{\scriptscriptstyle\text{Ne}}q$ is the dispersion relation (in the long wavelength limit) that only depends on the magnitude (not direction) of the wavevector, $q=|\bm{q}|=\sqrt{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}}$, $c_{\scriptscriptstyle\text{Ne}}=1.133\times 10^{5}$ cm s-1 is the longitudinal sound velocity [32]; $\bm{q}/q$ is the unit vector in the longitudinal wavefront direction.
@@
$\bm{v}(\bm{r},t)=\sum_{\bm{q}}\frac{\bm{q}}{{\rm i}q}\sqrt{\frac{\hbar\omega_{q}}{2V\rho_{\scriptscriptstyle\text{Ne}}}}(a_{\bm{q}}{\rm e}^{-{\rm i}\omega_{q}t}+a^{\dagger}_{-\bm{q}}{\rm e}^{+{\rm i}\omega_{q}t}){\rm e}^{{\rm i}\bm{q}\cdot\bm{r}},$ (14)
and the acoustic velocity potential $\mathit{\Phi}(\bm{r},t)$ is
@@
$\partial_{t}\varrho(\bm{r},t)+\rho_{\scriptscriptstyle\text{Ne}}\nabla\cdot\bm{v}(\bm{r},t)=0.$ (18)
In our system, phonons are confined in a semi-infinite space, $z<0$, and the system is translationally invariant in xyx-y plane. We define the 2D lateral positions $\bm{r}_{\scriptscriptstyle\parallel}=(x,y)$ and wavevectors $\bm{q}_{\scriptscriptstyle\parallel}=(q_{x},q_{y})$. The density variation field in the Schrödinger picture $(t=0)$ is
@@
$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}),$ (19)
which only includes the modes when the density variation vanishes at the Ne-vacuum interface, $z=0$, where pressure is zero. The summation in Eq. (19) only includes $q_{z}>0$ to ensure that the modes are counted correctly. Likewise, the acoustic velocity potential is
@@
$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}+\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}),$ (20)
@@
$\times(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (21)
@@
The total Hamiltonian consists of the electron part $\hat{\mathcal{H}}_{\text{el}}$ as given above, the phonon part $\hat{\mathcal{H}}_{\text{ph}}$, and the electron-phonon coupling part $\hat{\mathcal{H}}_{\text{el-ph}}$,
$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\text{el}}+\hat{\mathcal{H}}_{\text{ph}}+\hat{\mathcal{H}}_{\text{el-ph}},$ (22)
@@
$\hat{\mathcal{H}}_{\text{ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a^{\dagger}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}a_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}},$ (23)
@@
$\hat{\mathcal{H}}_{\text{el-ph}}=\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z){\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}}(\hat{a}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hat{a}^{\dagger}_{-\bm{q}_{\scriptscriptstyle\parallel},q_{z}}).$ (24)
@@
The relaxation rate corresponding to electron scattered by a single phonon can be calculated from the Fermi golden rule, we write the electron-phonon coupling Hamiltonian via the modulation of dielectric constant as $\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$ and the displacement of neon surface as $\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{dis})}$, then the relaxation rate corresponding to each mechanisms and different thermal phonons can be expressed as
@@
$\times\delta(\sum_{q}\hbar\omega_{q}n_{q}-\sum_{q}\hbar\omega_{q}n^{\prime}_{q}-\hbar\omega_{x}),$
@@
$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (25)
$\mathit{\Gamma}^{(\text{mod})}$ $=\frac{2\pi}{\hbar}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}\left|\mathcal{V}^{\text{(dis)}}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}(\bm{r}_{\scriptscriptstyle\parallel},z)\right|_{10}^{2}G_{\bm{q}_{\scriptscriptstyle\parallel}}$
$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (26)
where the subscript “10” denotes the excited to ground state transition and $n_{q}=1/[\exp(\hbar\omega_{q}/(k_{\scriptscriptstyle\text{B}}T))-1]\ll 1$ denotes the average thermal occupation number of the phonon state with the wavenumber $q$ at $T=10$ mK. The function $G_{\bm{q}_{\scriptscriptstyle\parallel}}$ is
@@
The phonon-induced dephasing ($\Gamma_{\phi}$) of electrons was estimated following the treatment used in ref [34] based on the ensemble-averaged mean-square phase increment:
$\overline{\langle[\phi_{10}(t)-\phi_{10}(0)]^{2}\rangle}$ $=\frac{1}{\hbar^{2}}\int_{0}^{t}dt^{\prime}\int_{0}^{t}dt^{\prime\prime}\overline{\langle E_{10}(t^{\prime})E_{10}(t^{\prime\prime})\rangle}$ (28)
$=\int_{0}^{t}dt^{\prime}\int_{0}^{t}dt^{\prime\prime}\Gamma_{\phi}\delta(t^{\prime}-t^{\prime\prime})=\Gamma_{\phi}t,$
where $\overline{\langle\cdot\rangle}$ is the average over the thermal states of phonons and
@@
is the energy difference between two electronic states. The energy correlator can be approximated by a $\delta$-function which gives the depahsing rate $\Gamma_{\phi}$. The average value of the energy difference Eq. (29) is given by
$\overline{\langle E_{10}(t^{\prime})E_{10}(t^{\prime\prime})\rangle}$ (30)
@@
$\times[(n_{q}+1)e^{-i\omega_{q}(t^{\prime}-t^{\prime\prime})}-n_{q}e^{i\omega_{q}(t^{\prime}-t^{\prime\prime})}],$
and the integral over time can be found as
$\int_{0}^{t}dt^{\prime}\int_{0}^{t}dt^{\prime\prime}e^{-i\omega_{q}(t^{\prime}-t^{\prime\prime})}=\frac{4\sin^{2}(\omega_{q}t/2)}{\omega^{2}_{q}}$ (31)
@@
$\mathit{\Gamma}_{\phi}=\frac{2\pi}{\hbar^{2}c_{\text{Ne}}}\sum_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}[$ $(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{11}$- (32)
$-(\mathcal{V}_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}>0}{\rm e}^{{\rm i}\bm{q}_{\scriptscriptstyle\parallel}\cdot\bm{r}_{\scriptscriptstyle\parallel}})_{00}]^{2}\delta(q).$
For one-phonon processes, the summation effectively vanishes due to the presence of $\delta(q)$, indicating that pure dephasing is negligible. Consequently, the decoherence time is governed by the relaxation process. In the following section, we focus on calculating the relaxation times for two distinct interaction mechanisms.